OpenStudy (anonymous):

a^4a^5 simplify the expression?

OpenStudy (curry):

A^4(1*A)

OpenStudy (curry):

a^4(1*a)

OpenStudy (curry):

u get it?

OpenStudy (anonymous):

no

OpenStudy (curry):

well multiply it out a^4 * a^5

OpenStudy (anonymous):

oh yeah i get now

jimthompson5910 (jim_thompson5910):

\[\large a^4a^5=a^{4+5}=a^9\] So \[\large a^4a^5=a^9\]